OpenStudy (anonymous):

Find the indicated term of the arithmetic sequence. The sequence is -4, -7, -10… Find term a41

OpenStudy (anonymous):

a=first term = -4 d= common difference = -3 Tn = a +(n-1)d = -4 +(n-1)(-3) Tn = -4 -3n +3 = -1 -3n

OpenStudy (anonymous):

then we want the 41st term so sub n=41

OpenStudy (dumbcow):

what he said side note: elecengineer, you might try posting tutoring ad on mathlessons.com, its worked for me

OpenStudy (anonymous):

T(41) = -1 -3(41) =-124

OpenStudy (anonymous):

so thats the answer -124?